OpenStudy (anonymous):

A girl flies a kite at a height of 300 ft, the wind carrying the kite horizontally away from her at a rate of 4 ft/sec. How fast must she let out the string when the kite is 428 ft away from her? On a problem like this would I differentiate a^2+b^2=c^2 or a^2+300=c^2? Because the 300 seems to be a constant. Just making sure for my test.

OpenStudy (anonymous):

a^2+300^2=c^2? ***

OpenStudy (irishboy123):

or a^2+300=c^2 careful, square the 300 too.

OpenStudy (anonymous):

Yeah caught that too hahah